### v1

# Nonparametric Estimation of Matching Efficiency and Mismatch in Labor Markets via Public Employment Security Offices in Japan, 1972-2024

Abstract
I examine changes in matching efficiency and elasticities in Japan’s labor market via Hello Work for unemployed workers from January 1972 to April 2024 using a nonparametric identification approach by Lange and Papageorgiou (2020). I find a declining trend in matching efficiency, consistent with decreasing job and worker finding rates. The implied match elasticity with respect to unemployment is 0.5-0.9, whereas the implied match elasticity with respect to vacancies is 0.1-0.4. Decomposing aggregate data into full-time and part-time ones, I find that the sharp decline of matching efficiency after 2015 shown in the aggregate trend is driven by the decline of both full-time and part-time ones. Second, I extend the mismatch index proposed by Şahin et al. (2014) to the nonparametric version and develop the computational methodology. I find that the mismatch across job categories is more severe than across prefectures and the original Cobb-Douglas mismatch index is underestimated.
Keywords: matching efficiency, mismatch
JEL code: E24, J61, J62, J64
1 Introduction
Understanding labor market matching efficiency and mismatch, which is defined as the discrepancy between the distribution of vacant jobs and unemployed workers, is crucial for comprehending labor market dynamics. However, few studies have investigated the contributions of mismatch to unemployment dynamics within the Japanese labor market. This paper seeks to address this gap in the literature by quantifying matching efficiency via the novel nonparametric approach of Lange and Papageorgiou (2020) and labor mismatch by extending the mismatch index proposed by Şahin et al. (2014) to a nonparametric version.
The matching function, which maps the number of hirings to the number of job seekers and recruiters, incorporating matching efficiency that includes their search effort, plays a central role in the literature on frictional labor markets. A significant issue identified in the extensive literature on estimating the matching function, as pointed out by Lange and Papageorgiou (2020), is the reliance on strong functional form assumptions, such as the Cobb-Douglas, which lack both empirical and theoretical grounding. If the Cobb-Douglas specification is overly restrictive, it can lead to biased estimates of matching efficiency and elasticity. To address this problem, Lange and Papageorgiou (2020) relax the functional form restrictions on the matching function, opting instead to non-parametrically identify the matching function, thereby recovering matching efficiency. To the best of my knowledge, this paper is the first to apply the nonparametric approach to the Japanese labor market.
Relatedly, several papers quantified various dimensions of mismatch and their contributions to the changes in the unemployment rates using the Cobb-Douglas specification. Şahin et al. (2014) develop a methodology to measure mismatch in various dimensions as the deviations of observed labor market outcomes from the socially efficient outcomes. In the paper, the Cobb-Douglas specification significantly simplifies the mismatch index by canceling out time-varying matching efficiency and elasticities. As a methodological contribution, I combine the nonparametric approach with the mismatch index allowing for nonparametrically recovered matching efficiency and elasticities and provide how to compute the index using Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012).
The main empirical findings of this paper are twofold. First, matching efficiency (normalized to 1972) in the labor market via Hello Work shows a declining trend with notable fluctuations, which is consistent with downward trends of job and worker finding rates. This might be because matching opportunity out of the government-operated platform has increased. Implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (the range in 0.5-0.7) and Japanese studies. On the other hand, implied match elasticity with respect to vacancies is 0.1-0.4, which is comparable to Lange and Papageorgiou (2020) (0.15-0.3) and Japanese studies.
Second, substantial mismatch are observed in both prefecture-level and job-category-level and the trend is upward with fluctuations. The mismatch across prefectures is 0.3 on average, whereas the mismatch across job categories is 0.4-0.65, larger than previous studies such as Shibata (2020) and Higashi and Sasaki (2023). The nonparametric mismatch index highlights the downward bias of the original Cobb-Douglas mismatch index through Monte Carlo simulation and empirical application.
1.1 Related Literature
This paper contributes to the three strands of literature. First, I examine the trend of matching efficiency in Japanese labor markets via Hello Work nonparametrically using a novel approach (Lange and Papageorgiou 2020), which shows how to nonparametrically identify the matching function and estimate the matching function allowing for unobserved matching efficacy, without imposing the usual independence assumption between matching efficiency and search on either side of the labor market, allowing for multiple types of jobseekers. Lange and Papageorgiou (2020) highlight positive correlations between efficiency and market structure such as tightness and so on, which induces a positive bias in the estimates of the vacancy elasticity whenever unobserved matching efficacy is not controlled for, as is the case in the traditional Cobb-Douglas matching function with constant elasticity parameters. [^1] Brancaccio et al. (2020a, 2023) apply the method to estimate matching function in exporter-ship transactions in the global bulk shipping market. Brancaccio et al. (2020b) summarize practical issues. Implementing their approach, I add updated results from the more flexible approach to the existing findings reported in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018) using the traditional Cobb-Douglas matching function with geographical and occupational category fixed effects to capture geographical and occupational heterogeneity. Table 1 summarizes the previous findings. My findings are also useful for comparison with other countries’ results reported in Bernstein et al. (2022) and Petrongolo and Pissarides (2001).
Table 1: Estimation results of the aggregate matching function in Japan
Paper CRS FE Estimation Sample $\frac{d\log M}{dlogU}$ $\frac{d\log M}{dlogV}$
Kano and Ohta (2005) yes yes OLS 1973–1999 0.56-0.59 0.29-0.3
Kambayashi and Ueno (2006) no yes OLS 1996–2001 0.81 0.30
Sasaki (2008) yes no OLS 1982–2016 0.38-0.45
Shibata (2013) yes no GMM 2000–2013 0.4
Shibata (2020) yes no GMM 2000–2019 0.34-0.4
Kawata et al. (2016) yes yes IV 2000–2009 0.51-0.6
Kawata (2019) yes yes OLS 2012–2016 0.52
Higashi (2018) yes no OLS 2008–2016 0.37-0.46 0.17-0.24
Higashi (2020) yes no OLS 2006–2016 0.48 0.17
Higashi (2021) yes no OLS, 2SLS 2008–2018 0.22-0.46 0.15-0.31
Higashi and Sasaki (2023) yes yes OLS 2016–2021 Calibrate Calibrate
Note: All models use a Cobb–Douglas matching function. Constant-returns-to-scale (CRS) imposes restriction on matching elasticity parameters in estimation. Fixed effect (FE) captures geographical-level and time-level heterogeneity. Estimation Instrument Variable (IV), Two-Stage-Least-Square (2SLS), and Generalized Method of Moments (GMM) based on Borowczyk-Martins et al. (2013) use instruments in the specific context. Blank cells mean no reported results.
Second, this paper contributes to the literature on mismatch Şahin et al. (2014). The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al. (2016). In the context of Japanese labor markets, several papers estimate mismatch at geographical- and job-category-level using Cobb-Douglas specification (Shibata 2013, 2020, Kawata et al. 2016, Higashi 2018, Kawata 2019, Higashi 2020, 2021, Higashi and Sasaki 2023). This paper points out that Cobb-Douglas specification cancels out time variations of matching efficiency which induce bias of the mismatch index. Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012), Dubé et al. (2012) to calculate the geographical- and job-category-level mismatch index.
Third, this paper contributes to the literature on the impact of COVID-19 on the matching efficiency of the worker-vacancy platform operated by the government. Due to the extensive volume of literature relevant to this topic, I will not provide a comprehensive list of all sources to maintain clarity and conciseness. The most closely related papers in my context are Higashi and Sasaki (2023) and Fukai et al. (2021). Higashi and Sasaki (2023) estimate mismatch indices for local labor markets clustered in by occupations vulnerable and not vulnerable to COVID-19 by assuming Cobb-Douglas matching function with calibrated matching elasticity parameters. Fukai et al. (2021) utilize the Labor Force Survey, a large-scale government statistic, to estimate the group average treatment effect of COVID-19 on employment status for each month from January to June 2020. To the best of my knowledge, this is the first paper to describe changes in matching efficiency before and after COVID-19 using a novel nonparametric approach.
2 Model
2.1 Nonparametric aggregate matching function
Our main interest is in matching efficiency and matching elasticity with respect to the number of unemployed workers and vacancies in the labor market via Public Employment Security Offices in Japan. A matching function based on search models plays a central role in labor economics. [^2] See Pissarides (2000), Petrongolo and Pissarides (2001), and Rogerson et al. (2005) for reference. The matching function relies on random search from both sides of the market, that is, individuals seeking jobs represent the supply of labor and recruiters represent the demand for labor. To estimate the matching function and recover matching efficiency, I follow the novel approach proposed by Lange and Papageorgiou (2020). [^3] Lange and Papageorgiou (2020) additionally incorporate search effort (Mukoyama et al. 2018) and recruitment index (Davis et al. 2013). Unfortunately, our Hello Work data does not report the information. The paper points out the endogeneity problem of matching efficiency (Borowczyk-Martins et al. 2013) and the problem of too restrictive specification of a Cobb-Douglas matching function with the fixed matching elasticity then proposes nonparametric identification and estimation of matching efficiency under some conditions introduced later.
Let unscripted capital letters $(A,U,V)$ denote random variables while realizations are subscripted by time $t$. I consider the matching function $m_{t}(\cdot,\cdot)$ that maps period-$t$ unemployed workers $U_{t}$, per-capita search efficacy/matching efficiency of the unemployed workers $A_{t}$, and vacancies $V_{t}$ into hires $H_{t}$. I assume that the underlying data generating process is stationary and that I observe a long enough time-series so that I can treat the joint distribution $G:\mathbb{R}_{+}^{3}\rightarrow[0,1]$ of $\left(H_{t},U_{t},V_{t}\right)$ as observed. Also, denote by $F(A,U)$ the joint distribution of $A$ and $U$.
I identify the matching function as well as unobserved, time-varying matching efficiency, $A.$ First, I assume that $V$ and $A$ are independent conditional on $U$, that is, $A\perp V\mid U$. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ has constant returns to scale (CRS). [^4] To follow the original model of Matzkin (2003), we can rewrite $H=m(AU,V)$ into $H/U=m(A,V/U)$ under CRS. Here, $H/U$ and $V/U$ are known as a job-finding rate and market tightness. Then, by applying nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) proves that $G(H,U,V)$ identifies $F(A,U)$ and $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point denoted as $A_{0}$ of the support of $(A,U,V)$. [^5] In Appendix, I report finite sample performance and methodological extension with Monte Carlo simulation. Based on simulation results with sample size $T=50$, the sample size in this paper is enough to recover matching efficiency well. The code is available on the author’s Github. Also, see Brancaccio et al. (2020b) for practical issues. The approach is used to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023).
2.2 Nonparametric mismatch
I construct a nonparametric version of a mismatch index which is a measure of the fraction of hires foregone due to mismatch. Suppose there are $G$ local markets in time $t$ with a given number of vacancies, unemployed workers, and aggregate matching efficiencies in the economy. The number of hires $H_{gt}$ in market $g$ in time $t$ is determined by a matching function $m(A_{gt}U_{gt},V_{gt})$. Note that, unlike Şahin et al. (2014), I allow any specification of $m$ satisfying CRS and independence of $A_{gt}$ and $U_{gt}$ conditional on $V_{gt}$. Given matching efficiency $A_{gt}$ and the number of vacancies $V_{gt}$, a social planner maximizes the total hires in the economy by distributing a given number of unemployed workers to each labor market.
Following Şahin et al. (2014), I define mismatch as the deviation of the hires in the data from the efficient allocation chosen by the social planner. Under the assumptions of homogenous productivity and job separation rate across labor markets, the optimal allocation of unemployed workers satisfies the following equilibrium conditions:
$\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)=\cdots=\frac{\partial m}{\partial U}\left(A_{Gt}U_{Gt}^{*},V_{Gt}\right)$ (1)
where * denotes the planner’s allocation. The equilibrium condition means that keeping total unemployed workers $U_{t}=\sum_{g=1}^{G}U_{gt}$, the planner allocates more unemployed workers to more effective labor markets, that is, with more vacancies and higher matching efficiency until their marginal contribution to the hires is equalized across markets. The main difference from Şahin et al. (2014) is that equilibrium condition (1) does not have an analytical formula like Cobb-Douglas specification. In Section 3.2, I explain how to compute the optimal allocation of unemployed workers numerically.
Using optimal allocation of unemployed workers $U_{gt}^{*}$, the aggregate actual and optimal number of new hires can be expressed as
$H_{t}$ $=\sum_{g=1}^{G}m(A_{gt}U_{gt},V_{gt}),$
$H_{t}^{*}$ $=\sum_{g=1}^{G}m(A_{gt}U_{gt}^{*},V_{gt}).$
Using these expressions, I define mismatch index as
$\mathcal{M}_{t}=1-\frac{H_{t}}{H_{t}^{*}}$ (2)
where $\mathcal{M}_{t}$ measures the fraction of hires lost in period $t$ because of misallocation. This index accounts for the heterogeneity of matching efficiencies across labor markets.
3 Estimation
3.1 Matching efficiency and elasticities
Following Lange and Papageorgiou (2020), I begin by estimating $F(A_{0}|U)$ across the support of $U$. To this end, we utilize the distribution of hires conditional on unemployed, $U$, and observed vacancies, $V$. Specifically, we have
$F(A_{0}|\psi U_{0})$ $=G_{H|U,V}(\psi H_{0}|\psi U_{0},\psi V_{0})\quad\text{for any arbitrary scalar }\psi.$
$F(\psi A_{0}|\lambda U_{0})$ $=G_{H|U,V}(\psi H_{0}|\lambda U_{0},\psi V_{0})\quad\text{where }\lambda>0\text{ is a scaling factor}$
where $F(A_{0}|\psi U_{0})$ and $G_{H|U,V}$ are conditional distributions. By varying $(\psi,\lambda)$, we can therefore trace out $F(A|U)$ across the entire support of $(A,U)$.
Given our finite data, we rely on an estimate of $G_{H|U,V}$ for our constructive estimator. Consider an arbitrary point $(H_{\tau},U_{\tau},V_{\tau})$. To obtain $G(H_{\tau}|U_{\tau},V_{\tau})$, we compute the proportion of observations with less than $H_{\tau}$ observed hires among observations proximate to $(U_{\tau},V_{\tau})$ in $(U,V)$-space. Practically, this is achieved by averaging across all observations in the data, penalizing those with values $(U_{t},V_{t})$ using a kernel that weighs down observations distant from $(U_{\tau},V_{\tau})$. Consequently, our estimate is given by
$F(\psi A_{0}|\lambda U_{0})=G_{H|U,V}(\psi H_{0}|\lambda U_{0},\psi V_{0})$
$\hat{F}(\psi A_{0}|\lambda U_{0})=\sum 1(H_{t}<\psi H_{0})\kappa(U_{t},V_{t},\lambda U_{0},\psi V_{0})$
where $\kappa(.)$ denotes a bivariate normal kernel with bandwidth 0.01.
Having recovered the distribution function $F(A|U)$, we invert $F(A_{t}|U_{t})$ to derive $A_{t}$. This is achieved by
$A_{t}=F^{-1}(G(H_{t}|U_{t},V_{t})|U_{t})$
for all observations $(H_{t},U_{t},V_{t})$ in the dataset. Finally, we recover the matching function as
$m(A_{t},U_{t})=G^{-1}(F(A_{t}|U_{t})|U_{t}).$
Finally, for calculating matching elasticities, I run a LASSO regression projecting hires on the original and squared numbers of vacancies and unemployed interacted with implied matching efficiency. The estimates approximate the derivatives of the matching function with respect to vacancies and unemployed interacted with implied matching efficiency, that is, an estimate of the elasticity of the matching function. [^6] The matching elasticity with respect to unemployed $\frac{d\log m(AU,V)}{d\log U}=\frac{dm(AU,V)}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{dAU}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{AU}{H}=\frac{d\log m(AU,V)}{d\log AU}$ is obtained from the regression coefficient of $H$ on $AU$ and multiplying it by $\frac{AU}{H}$. Concretely, we approximate $m$ by the second order polynomial $m=\beta_{1}(AU)+\beta_{2}(AU)V+\beta_{3}V+\beta_{4}(AU)^{2}+\beta_{5}V^{2}$ and get $\frac{dm(AU,V)}{dAU}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$ and $\frac{d\log m}{d\log U}=\frac{d\log m}{d\log AU}=(\beta_{1}+\beta_{2}V+2\beta_{4}(AU))\frac{U}{H}$ and $\frac{d\log m}{d\log V}=(\beta_{2}(AU)+\beta_{3}+2\beta_{5}V)\frac{V}{H}$. For calculating mismatch, we need $\frac{dm}{dU}=\frac{d\log m(AU,V)}{d\log AU}\frac{H}{U}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$.
3.2 Mismatch
Unlike the original Cobb-Douglas specification of Şahin et al. (2014), equilibrium condition (1) does not have an analytical formula. Instead, I employ the Mathematical Programming with Equilibrium Constraint (MPEC) approach proposed by Su and Judd (2012) and Dubé et al. (2012). The constrained optimization problem in time $t$ is defined as follows.
$\max_{U_{1t},\cdots,U_{Gt}}$ $\sum_{g=1}^{G}m_{gt}\left(A_{gt}U_{gt},V_{gt}\right),$ (3)
subject to $\text{subject to }\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)$ $=\cdots=\frac{\partial m}{\partial U}\left(A_{Gt}U_{Gt}^{*},V_{Gt}\right),$
$U_{t}$ $=\sum_{g=1}^{G}U_{gt}.$
Maximizing the objective function is equivalent to solving equilibrium conditions (1) so that the central planner’s problem represents a nonlinear system of equations with linear equilibrium constraints. I solve the model by using Ipopt.jl and JuMP.jl which is often used for the MPEC approach. Note that the formulation allows us to incorporate a rich set of additional equilibrium constraints and elements such as the government’s costs of moving unemployed workers from some markets to other markets for calibration studies, which is out of the scope of this paper.
4 Data
First, I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for the month-level aggregate data from January 1972 to April 2024. These datasets include the number of job openings, job seekers, and successful job placements, primarily sourced from the Ministry of Health, Labour and Welfare (MHLW) of Japan, which publishes monthly reports and statistical data on the Public Employment Security Office, commonly known as Hello Work. Hello Work is a government-operated institution in Japan that provides job seekers with employment counseling, job placement services, and vocational training, playing a critical role in Japan’s labor market. The data is often used as in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018) estimating the traditional Cobb-Douglas matching function. Using up-to-date Hello Work data, Kawata and Sato (2021) construct a simple framework to quantitatively measure the impacts of an economic shock of COVID-19 on unemployed workers’ welfare in their companion project. [^7] https://www.crepe.e.u-tokyo.ac.jp/material/crepecl12.html: Accessed 2024 June 6. The period for my dataset is selected to ensure the longest consistent timeframe available at the time of writing this paper. In Appendix A, we provide additional analysis using the year-level aggregate data in more extended periods, available from January 1963 to April 2024, instead of month-level one.
The three types of number of job openings, job seekers, and successful job placements are reported in the data. The first is the number for full-time and part-time workers and jobs. The second is the number for full-time workers and jobs. The third is the number for part-time workers and jobs. The first is the sum of the second and third. Using the three types of data, I can decompose the labor market features into full-time and part-time ones.
Second, for estimating nonparametrically worker-job mismatch across regions and occupation categories, I use these submarket data in 2012-2024. The period limitation shared with Kawata (2019) and Higashi (2018) is due to the revision of job classifications before and after 2012, which makes it difficult to accurately connect the data. Finally, I use all 47 prefectures and 67 job categories in each month as submarkets.
5 Empirical Results
We apply the above estimation approach for each dataset. Before presenting the results, I test the assumption that vacancies are independent of matching efficiency, conditional on overall labor supply, i.e., $V\perp A\mid U$. Specifically, we use the residuals from a regression of vacancies $V$ on the unemployed $U$, and similarly, the residuals of implied matching efficiency $A$ on $S$. For the aggregate data, the correlation between these two residuals is close to zero (-0.04), indicating no systematic relationship between them. Conversely, for the full-time and part-time data, the correlations between the residuals are 0.12 and -0.14. Because the patterns appear to be influenced by some outliers as in Figure 1, the effect of violation of independence seems limited.
Figure 1: Residual plot
5.1 Aggregate trends in 1972-2023
Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($M/U$ and $M/V$) in aggregate labor markets. In summary, the numbers of unemployed individuals and vacancies increase with fluctuations, corresponding with market tightness, while the number of hires shows a noticeable decline around the late 1970s, peaks and troughs from the mid-1980s to the late 1990s, peaks around mid-2000s and 2010, and a sharp decline towards recent years. Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the previous studies (Elsby et al. 2015).
Figures 2 (e) and (f) present the estimation results of the matching function along with matching efficiency and elasticities ($\frac{d\ln M}{d\ln AU}$ and $\frac{d\ln M}{d\ln V}$). Notably, matching efficiency (normalized to 1972) shows a declining trend with notable fluctuations, which is consistent with the downward trends of job and worker finding rates. In particular, the significant decline after 2015 is remarkable. This seems due to an increase in matching opportunities outside of the government-operated platform.
The implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (range: 0.5-0.7) and Japanese studies such as Higashi (2018) (0.38 for 2000-2014 monthly), Kawata (2019) (0.48 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.56 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.6 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.8 for 1996-2001 prefecture-month-level). [^8] For reference, Petrongolo and Pissarides (2001) summarize early aggregate studies in many countries based on a Cobb-Douglas matching function with the flow of hires on the left-hand side and the stock of unemployment and job vacancies on the right-hand side. In short, match elasticity with respect to unemployment is in the range of 0.5–0.7. Bernstein et al. (2022) review the recent empirical literature that estimates the matching elasticity in the U.S. by four types of specification; Cobb-Douglas, Cobb-Douglas with endogeneity correction, a constant elasticity of substitution (CES), and nonparametric (Lange and Papageorgiou 2020). The broad range of estimates is partly due to variations in data choices and periods. Higher estimates are typically derived from JOLTS data (Borowczyk-Martins et al. 2013, Şahin et al. 2014), while lower estimates are often obtained from CPS flows data (Barnichon and Figura 2015) or job-category-level hires from CPS data (Şahin et al. 2014). On the other hand, the implied match elasticity with respect to vacancies is 0.1-0.4, which is comparable to Lange and Papageorgiou (2020) (range: 0.15-0.3) and Japanese studies such as Higashi (2018) (0.24 for 2000-2014 monthly), Kawata (2019) (0.52 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.3 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.2 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.3 for 1996-2001 prefecture-month-level).
Figures 2 (g) and (h) illustrate some correlation patterns between matching efficiency and market structure variables such as labor market tightness, worker finding rate, and job finding rate. Consistent with Lange and Papageorgiou (2020), these highlight positive correlations between efficiency and market structure, such as tightness, which induce a positive bias in the estimates of the vacancy elasticity whenever unobserved matching efficiency is not controlled for, as is the case in traditional estimators.
Figure 2: Month-level aggregate results 1972-2024
5.2 Decomposition of part-time and full time trends in 1972-2023
Next, we decompose the aggregate trends into full-time and part-time labor markets’ trends. Figure 3 provide full-time and part-time labor markets’ trends corresponding Figures 2 (a)-(f). Panel (a) shows that full-time unemployment and vacancies show higher counts and more significant fluctuations compared to part-time data and that part-time unemployment and vacancies exhibit a more gradual and steady trend over the years. Both full-time and part-time labor markets experience periods of high tightness during the same general periods, notably in the early 1990s. Part-time labor market tightness shows more pronounced peaks compared to the full-time labor market. Panel (b) shows that the sharp decline in full-time hires post-2008 suggests greater sensitivity to economic downturns and structural changes in the labor market. The steady rise in part-time hires indicates increasing reliance on part-time employment, possibly due to greater flexibility and adaptability in the labor market. Panel (c) illustrates that full-time employment data covers a wider range of both unemployed and vacancies, indicating higher variability in the full-time labor market in contrast to a more stable part-time labor market. Panel (d) shows that full-time finding rates show higher initial values and more significant declines compared to part-time finding rates. Part-time finding rates demonstrate a more gradual decline and stable fluctuations over time. Full-time finding rates exhibit larger fluctuations and more pronounced declines, indicating greater sensitivity to economic changes. Part-time finding rates, while also showing initial declines, stabilize earlier and display less volatility, suggesting a more resilient part-time labor market.
Given estimation results in each dataset, Figure 4 (e) depicts the trend of implied matching efficiency. The downward trends are common. Full-time matching efficiency fluctuates from 80% to 300%, whereas part-time matching efficiency remains stable at 60% compared to January 1972. This concludes that the sharp decline of matching efficiency after COVID-19 shown in the aggregate trend is driven by the decline of full-time matching efficiency.
Figure 4 (f) depicts the trends of matching elasticities. For matching elasticity with respect to unemployed, the full-time trend exhibits significant fluctuations, especially before the 1980s, ranging from 0.2 to around 1.0, indicating varying sensitivity of matching efficiency to changes in unemployment over time. In contrast, the part-time trend peaks, ranging from 0.9 to 1.0 after 2020, which may indicate the post-COVID effect on the elasticity. For matching elasticity with respect to vacancies, the full-time trend shows a relatively stable trend, starting around 0.2, with slight increases in recent years, whereas the part-time trend maintains lower and more stable values around 0.1 to 0.3, suggesting that vacancies have a less pronounced effect on matching efficiency in the part-time labor market.
Figure 4 illustrates some correlation patterns between matching efficiency and market structure variables corresponding with Figures 2 (g)-(h). I find the stronger and more consistent relationships in the full-time labor market. This confirms that positive correlations between efficiency and market structure driven mainly by full-time markets.
Figure 3: Month-level full-time and part-time results 1972-2024
Figure 4: Month-level full-time and part-time results 1972-2024 (Continued)
5.3 Prefecture-level aggregate results in 2012-2024
Figure 5 displays the month-prefecture level matching efficiency, normalized to 2013 January Tokyo, across different regions of Japan from 2012 to 2024. Each panel shows the efficiency trends within a specific region, highlighting notable regional variations. For instance, the Tohoku region (panel b) and Chubu region (panel d), which are rural areas, exhibit relatively high volatility and generally higher efficiency values compared to other regions, suggesting greater variability in matching efficiency. In contrast, regions such as Kanto (panel c) and Kansai (panel e) show more stable and lower efficiency levels. The Tohoku area is about twice as efficient as the Tokyo area. These patterns suggest that regional economic factors and labor market conditions significantly influence matching efficiency across Japan.
Figure 5: Month-prefecture level matching efficiency 2012-2024
Figures 6 and 7 illustrate month-prefecture-level matching elasticities with respect to unemployed and vacancies in 2012-2024. There is significant regional heterogeneity within the specific region. The matching elasticities with respect to unemployed and vacancies are 0.6-0.8 and 0.05-0.3 on average.
Figure 6: Month-prefecture level matching elasticity with respect to unemployed 2012-2024
Figure 7: Month-prefecture level matching elasticity with respect to vacancies 2012-2024
5.4 Job category level results in 2012-2024
Figure 8 illustrates significant disparities in matching efficiency across job categories, normalized to 2013 January for clerical jobs. Primary and secondary industries show relatively higher efficiency than tertiary industries. For example, categories like ”Agriculture, Forestry, and Fishing” (panel f) and ”Construction and Mining” (panel i) in primary and secondary industries exhibit high volatility and generally higher efficiency levels. In contrast, job categories such as ”Managerial” (panel a) and ”Service” (panel e) display more stable and lower efficiency values. These variations highlight the diverse nature of labor market dynamics across different sectors, influenced by factors such as skill specialization, economic cycles, and industry-specific demand fluctuations.
Figure 8: Month-job category level matching efficiency results 2012-2024
Figures 9 and 10 illustrate matching elasticities with respect to unemployed and vacancies in each job category in 2012-2024. As in the prefecture-level results, there is significant regional heterogeneity within the specific job area. The matching elasticities with respect to unemployed and vacancies fluctuates with a range generally between 0.1 and 6.0, which is larger than prefecture-level results. This indicates that matching heterogeneity depends on the definition of markets and matters in job-category level rather than geographical level.
Figure 9: Month-job category level matching elasticity with respect to unemployed 2012-2024
Figure 10: Month-job category level matching elasticity with respect to vacancies 2012-2024
5.5 Mismatch in 2012-2024
Finally, I compute the estimated nonparametric mismatch index in 2012-2024. Panel (a) of Figure 11 illustrates the estimated nonparametric mismatch index across 47 prefectures from 2014 to 2024. The index, representing the deviation of the actual number of hires from a social planner’s optimal one, demonstrates significant seasonal fluctuations throughout the period. Mismatch across prefectures increases to around 0.3. Notably, the mismatch index shows a cyclical pattern with peaks and troughs, suggesting seasonal factors or regional economic cycles influencing labor market dynamics. The mismatch level across prefectures is larger than previous Cobb-Douglas specification results such as Shibata (2020) and Higashi and Sasaki (2023).
Panel (b) of Figure 11 shows the mismatch index across different job categories over a similar period. Unlike the prefecture-based mismatch, this index exhibits a more consistent upward trajectory, with less pronounced seasonal variations. The mismatch index across job categories increases to 0.6, which is significantly larger than Shibata (2020) and Higashi and Sasaki (2023), which indicates that the standard Cobb-Douglas specification underestimates the mismatch index due to the absence of time-varying matching efficiency and elasticity. The steady increase in the mismatch index across job categories points to a growing dissonance between the skills offered by job seekers and the skills demanded by employers. This trend may reflect evolving industry demands, technological advancements, and the lag in the workforce’s adaptation to new skills. The sharper rise in the mismatch index post-2020 further suggests that the pandemic may have exacerbated skill mismatches, possibly due to shifts in industry dynamics, such as the acceleration of digital transformation and the differing impacts on various sectors.
Figure 11: Estimated nonparametric mismatch index
6 Conclusion
I investigate how matching efficiency in the labor market via Public Employment Security Offices in Japan for unemployed workers changed during the period 1972-2024. By applying a novel nonparametric identification approach proposed by Lange and Papageorgiou (2020) to monthly data, I find that matching efficiency (normalized to 1972) exhibits a declining trend with notable fluctuations, consistent with the downward trends in job and worker finding rates. Finally, I extend the mismatch index proposed by Şahin et al. (2014) to the nonparametric version and find that the mismatch across job categories is more severe than across prefectures and the original Cobb-Douglas mismatch index is biased.
Acknowledgments
I thank Fuhito Kojima, Kosuke Uetake, Akira Matsushita, Kazuhiro Teramoto, Ryo Kambayashi, Higashi Yudai for their valuable advice. This work was supported by JST ERATO Grant Number JPMJER2301, Japan.
Appendix A Appendix
A.1 Year-level trends in 1966-2023
Figure 12: Year-level results 1966-2024
Figures 12 provide a year-level counterpart to Figure 2. The findings in the main text remain valid.

### v2

@@
 I find a declining trend in matching efficiency, consistent with decreasing job and worker finding rates.
-The implied match elasticity with respect to unemployment is 0.5-0.9, whereas the implied match elasticity with respect to vacancies is 0.1-0.4.
+The implied match elasticity with respect to unemployment is 0.5-0.9, whereas the implied match elasticity with respect to vacancies varies between -0.4 and 0.4.
 Decomposing aggregate data into full-time and part-time ones, I find that the sharp decline of matching efficiency after 2015 shown in the aggregate trend is driven by the decline of both full-time and part-time ones.
@@
 (2014) to the nonparametric version and develop the computational methodology.
-I find that the mismatch across job categories is more severe than across prefectures and the original Cobb-Douglas mismatch index is underestimated.
+I find that the mismatch across occupations is more severe than across prefectures and the original Cobb-Douglas mismatch index is underestimated.
 Keywords: matching efficiency, mismatch
@@
 Implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (the range in 0.5-0.7) and Japanese studies.
-On the other hand, implied match elasticity with respect to vacancies is 0.1-0.4, which is comparable to Lange and Papageorgiou (2020) (0.15-0.3) and Japanese studies.
-Second, substantial mismatch are observed in both prefecture-level and job-category-level and the trend is upward with fluctuations.
-The mismatch across prefectures is 0.3 on average, whereas the mismatch across job categories is 0.4-0.65, larger than previous studies such as Shibata (2020) and Higashi and Sasaki (2023).
+On the other hand, implied match elasticity with respect to vacancies varies between -0.4 and 0.4, which is more volatile than Lange and Papageorgiou (2020) (0.15-0.3) and Japanese studies (0.15-0.6).
+Second, substantial mismatch are observed in both prefecture-level and occupation-level and the trend is upward with fluctuations.
+The mismatch across prefectures is 0.3 on average, whereas the mismatch across occupations is 0.4-0.65, larger than previous studies such as Shibata (2020) and Higashi and Sasaki (2023).
 The nonparametric mismatch index highlights the downward bias of the original Cobb-Douglas mismatch index through Monte Carlo simulation and empirical application.
@@
 Blank cells mean no reported results.
-Second, this paper contributes to the literature on mismatch Şahin et al.
-(2014).
+Second, this paper contributes to the literature on mismatch (Şahin et al.
+2014).
 The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al.
 (2016).
-In the context of Japanese labor markets, several papers estimate mismatch at geographical- and job-category-level using Cobb-Douglas specification (Shibata 2013, 2020, Kawata et al.
-2016, Higashi 2018, Kawata 2019, Higashi 2020, 2021, Higashi and Sasaki 2023).
+In the context of Japanese labor markets, several papers estimate mismatch at geographical- and occupation-level using Cobb-Douglas specification (Shibata (2013, 2020), Kawata et al.
+(2016), Higashi (2018), Kawata (2019), Higashi (2020, 2021), Higashi and Sasaki (2023)).
 This paper points out that Cobb-Douglas specification cancels out time variations of matching efficiency which induce bias of the mismatch index.
-Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012), Dubé et al.
-(2012) to calculate the geographical- and job-category-level mismatch index.
+Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012) and Dubé et al.
+(2012) to calculate the geographical- and occupation-level mismatch index.
 Third, this paper contributes to the literature on the impact of COVID-19 on the matching efficiency of the worker-vacancy platform operated by the government.
@@
 Concretely, we approximate $m$ by the second order polynomial $m=\beta_{1}(AU)+\beta_{2}(AU)V+\beta_{3}V+\beta_{4}(AU)^{2}+\beta_{5}V^{2}$ and get $\frac{dm(AU,V)}{dAU}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$ and $\frac{d\log m}{d\log U}=\frac{d\log m}{d\log AU}=(\beta_{1}+\beta_{2}V+2\beta_{4}(AU))\frac{U}{H}$ and $\frac{d\log m}{d\log V}=(\beta_{2}(AU)+\beta_{3}+2\beta_{5}V)\frac{V}{H}$.
-For calculating mismatch, we need $\frac{dm}{dU}=\frac{d\log m(AU,V)}{d\log AU}\frac{H}{U}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$.
+For calculating mismatch, we need $\frac{dm}{dU}=\frac{d\log m(AU,V)}{d\log AU}\frac{H}{U}$ at $U=U^{*}$ and $H=H^{*}=m(AU^{*},V)$.
 3.2 Mismatch
@@
 Note that the formulation allows us to incorporate a rich set of additional equilibrium constraints and elements such as the government’s costs of moving unemployed workers from some markets to other markets for calibration studies, which is out of the scope of this paper.
+4 Monte Carlo simulation
+  • How large is the bias of matching efficiency from the Cobb Douglass specification?
+  • How large is the bias of mismatch from the Cobb Douglass specification?
+  • Mismatch literature:
+    • mismatch Shimer (2007)
+    • mismatch and large firms (Eeckhout 2018)
+    • Classical Jackman and Roper (1987)
 5 Data
@@
 In Appendix A, we provide additional analysis using the year-level aggregate data in more extended periods, available from January 1963 to April 2024, instead of month-level one.
+We study the labor markets for full-time and part-time workers and jobs.
+In the Hello Work system, part-time workers are defined as those who work fewer hours than regular employees at the same establishment.
+MHLW classifies part-time workers into two categories: (1) regular part-time workers, who have contracts for an indefinite period or a period exceeding four months, and (2) temporary part-time workers, whose contracts last between one and four months or whose employment is fixed and typically tied to seasonal demand.
+While in many countries, part-time and full-time jobs are distinguished by the number of hours worked, in Japan, these terms often imply additional differences, such as variations in responsibilities, benefits, and flexibility.
+Part-time workers generally have fewer responsibilities and benefits but enjoy more flexible working hours compared to full-time employees.
+Full-time workers are those who are not part-time workers.
 The three types of number of job openings, job seekers, and successful job placements are reported in the data.
@@
 The first is the sum of the second and third.
+Note that data on unemployed, vacancies, and hires exclude workers newly graduated from college, and are used to calculate the unemployment rate for Japan.
 Using the three types of data, I can decompose the labor market features into full-time and part-time ones.
-Second, for estimating nonparametrically worker-job mismatch across regions and occupation categories, I use these submarket data in 2012-2024.
-The period limitation shared with Kawata (2019) and Higashi (2018) is due to the revision of job classifications before and after 2012, which makes it difficult to accurately connect the data.
-Finally, I use all 47 prefectures and 67 job categories in each month as submarkets.
+Second, to estimate worker-job mismatch nonparametrically across regions and occupation categories, I use submarket data from January 2012 to March 2023 for prefecture-level analysis, and up to March 2024 for occupation category-level analysis.
+The time frame, consistent with Kawata (2019) and Higashi (2018), is constrained by the revision of job classifications before and after 2012, which complicates accurate data connection.
+Finally, I analyze all 47 prefectures and 67 occupations, treating each as a submarket on a monthly basis.
 6 Empirical Results
@@
 Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($M/U$ and $M/V$) in aggregate labor markets.
+[^8] Worker finding rates are also known as vacancy yields.
+Davis et al.
+(2013) define vacancy yields in month $t$ as the number of hires in month $t\left(h_{t}\right)$ divided by the number of vacancies in the previous month $\left(v_{t-1}\right)$ instead of the same month $\left(v_{t}\right)$, the two series are very similar.
 In summary, the numbers of unemployed individuals and vacancies increase with fluctuations, corresponding with market tightness, while the number of hires shows a noticeable decline around the late 1970s, peaks and troughs from the mid-1980s to the late 1990s, peaks around mid-2000s and 2010, and a sharp decline towards recent years.
-Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the previous studies (Elsby et al.
+Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the U.S.
+studies (Elsby et al.
 2015).
@@
 In particular, the significant decline after 2015 is remarkable.
-This seems due to an increase in matching opportunities outside of the government-operated platform.
+This seems due to an increase in matching opportunities outside of the government-operated platform, which is discussed in Otani (2024) using proprietary on-the-job search platform data.
 The implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (range: 0.5-0.7) and Japanese studies such as Higashi (2018) (0.38 for 2000-2014 monthly), Kawata (2019) (0.48 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.56 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.6 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.8 for 1996-2001 prefecture-month-level).
@@
 2013, Şahin et al.
-2014), while lower estimates are often obtained from CPS flows data (Barnichon and Figura 2015) or job-category-level hires from CPS data (Şahin et al.
+2014), while lower estimates are often obtained from CPS flows data (Barnichon and Figura 2015) or occupation-level hires from CPS data (Şahin et al.
 2014).
-On the other hand, the implied match elasticity with respect to vacancies is 0.1-0.4, which is comparable to Lange and Papageorgiou (2020) (range: 0.15-0.3) and Japanese studies such as Higashi (2018) (0.24 for 2000-2014 monthly), Kawata (2019) (0.52 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.3 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.2 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.3 for 1996-2001 prefecture-month-level).
+On the other hand, the implied match elasticity with respect to vacancies is 0.1-0.4 before 2015, which is comparable to Lange and Papageorgiou (2020) (range: 0.15-0.3) and Japanese studies such as Higashi (2018) (0.24 for 2000-2014 monthly), Kawata (2019) (0.52 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.3 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.2 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.3 for 1996-2001 prefecture-month-level).
+However, the elasticity declines significantly after 2010, reaching close to -0.5 by 2020.
+The negative elasticity implies that an increase in vacancies results in a decreasing number of matches, which could reflect structural changes in the labor market during COVID-19.
+The recent sharp decline also highlights a substantial inefficiency in how vacancies are being filled in the labor market.
 Figures 2 (g) and (h) illustrate some correlation patterns between matching efficiency and market structure variables such as labor market tightness, worker finding rate, and job finding rate.
@@
 Figure 2: Month-level aggregate results 1972-2024
-5.2 Decomposition of part-time and full time trends in 1972-2023
+6.2 Decomposition of part-time and full-time trends in 1972-2023
 Next, we decompose the aggregate trends into full-time and part-time labor markets’ trends.
@@
 The downward trends are common.
-Full-time matching efficiency fluctuates from 80% to 300%, whereas part-time matching efficiency remains stable at 60% compared to January 1972.
-This concludes that the sharp decline of matching efficiency after COVID-19 shown in the aggregate trend is driven by the decline of full-time matching efficiency.
+Full-time matching efficiency fluctuates from 80% to 300%, whereas part-time matching efficiency fluctuates from 60% to 300% compared to January 1972.
+This concludes that the sharp decline of matching efficiency after COVID-19 shown in the aggregate trend is driven by the decline of matching efficiency for both full-time and part-time workers.
 Figure 4 (f) depicts the trends of matching elasticities.
-For matching elasticity with respect to unemployed, the full-time trend exhibits significant fluctuations, especially before the 1980s, ranging from 0.2 to around 1.0, indicating varying sensitivity of matching efficiency to changes in unemployment over time.
-In contrast, the part-time trend peaks, ranging from 0.9 to 1.0 after 2020, which may indicate the post-COVID effect on the elasticity.
-For matching elasticity with respect to vacancies, the full-time trend shows a relatively stable trend, starting around 0.2, with slight increases in recent years, whereas the part-time trend maintains lower and more stable values around 0.1 to 0.3, suggesting that vacancies have a less pronounced effect on matching efficiency in the part-time labor market.
+The elasticity with respect to unemployment for full-time workers shows moderate fluctuations, particularly between the 1970s and the early 1980s, where it oscillates between 0.4 and 0.8.
+After the 1980s, this elasticity stabilizes around 0.5.
+In contrast, the part-time elasticity with respect to unemployment follows a more unstable pattern, fluctuating between 0.1 and 0.9 with significant variations over time.
+This instability highlights the vulnerability of part-time workers labor markets to economic shocks, particularly during periods of recession, such as the COVID-19 pandemic.
+The larger fluctuations in elasticity suggest that changes in unemployment among part-time workers have had an uneven and more pronounced impact on the matching process compared to full-time workers.
 Figure 4 illustrates some correlation patterns between matching efficiency and market structure variables corresponding with Figures 2 (g)-(h).
-I find the stronger and more consistent relationships in the full-time labor market.
+I find stronger and more consistent relationships in the full-time labor market.
 This confirms that positive correlations between efficiency and market structure driven mainly by full-time markets.
@@
 For instance, the Tohoku region (panel b) and Chubu region (panel d), which are rural areas, exhibit relatively high volatility and generally higher efficiency values compared to other regions, suggesting greater variability in matching efficiency.
-In contrast, regions such as Kanto (panel c) and Kansai (panel e) show more stable and lower efficiency levels.
+In contrast, urban regions such as Kanto (panel c) and Kansai (panel e) show more stable and lower efficiency levels.
 The Tohoku area is about twice as efficient as the Tokyo area.
@@
 Figure 7: Month-prefecture level matching elasticity with respect to vacancies 2012-2024
-5.4 Job category level results in 2012-2024
-Figure 8 illustrates significant disparities in matching efficiency across job categories, normalized to 2013 January for clerical jobs.
+6.4 Occupation level results in 2012-2024
+Figure 8 illustrates significant disparities in matching efficiency across occupations, normalized to 2013 January for clerical jobs.
 Primary and secondary industries show relatively higher efficiency than tertiary industries.
-For example, categories like ”Agriculture, Forestry, and Fishing” (panel f) and ”Construction and Mining” (panel i) in primary and secondary industries exhibit high volatility and generally higher efficiency levels.
-In contrast, job categories such as ”Managerial” (panel a) and ”Service” (panel e) display more stable and lower efficiency values.
+For example, categories like “Agriculture, Forestry, and Fishing” (panel f) and “Construction and Mining” (panel i) in primary and secondary industries exhibit high volatility and generally higher efficiency levels.
+In contrast, occupations such as “Managerial” (panel a) and “Service” (panel e) display more stable and lower efficiency values.
 These variations highlight the diverse nature of labor market dynamics across different sectors, influenced by factors such as skill specialization, economic cycles, and industry-specific demand fluctuations.
-Figure 8: Month-job category level matching efficiency results 2012-2024
-Figures 9 and 10 illustrate matching elasticities with respect to unemployed and vacancies in each job category in 2012-2024.
+Figure 8: Month-occupation level matching efficiency results 2012-2024
+Figures 9 and 10 illustrate matching elasticities with respect to unemployed and vacancies in each occupation in 2012-2024.
 As in the prefecture-level results, there is significant regional heterogeneity within the specific job area.
 The matching elasticities with respect to unemployed and vacancies fluctuates with a range generally between 0.1 and 6.0, which is larger than prefecture-level results.
-This indicates that matching heterogeneity depends on the definition of markets and matters in job-category level rather than geographical level.
-Figure 9: Month-job category level matching elasticity with respect to unemployed 2012-2024
-Figure 10: Month-job category level matching elasticity with respect to vacancies 2012-2024
+This indicates that matching heterogeneity depends on the definition of markets and matters in occupation level rather than geographical level.
+Figure 9: Month-occupation level matching elasticity with respect to unemployed 2012-2024
+Figure 10: Month-occupation level matching elasticity with respect to vacancies 2012-2024
 6.5 Mismatch in 2012-2024
-Finally, I compute the estimated nonparametric mismatch index in 2012-2024.
+Finally, I compute the estimated nonparametric mismatch index.
 Panel (a) of Figure 11 illustrates the estimated nonparametric mismatch index across 47 prefectures from 2014 to 2024.
@@
 Notably, the mismatch index shows a cyclical pattern with peaks and troughs, suggesting seasonal factors or regional economic cycles influencing labor market dynamics.
-The mismatch level across prefectures is larger than previous Cobb-Douglas specification results such as Shibata (2020) and Higashi and Sasaki (2023).
-Panel (b) of Figure 11 shows the mismatch index across different job categories over a similar period.
+The mismatch level across prefectures is larger than previous Cobb-Douglas specification results such as Shibata (2020) (0.05-0.1) and Higashi and Sasaki (2023) (0.15-0.25).
+Panel (b) of Figure 11 shows the mismatch index across different occupations over a similar period.
 Unlike the prefecture-based mismatch, this index exhibits a more consistent upward trajectory, with less pronounced seasonal variations.
-The mismatch index across job categories increases to 0.6, which is significantly larger than Shibata (2020) and Higashi and Sasaki (2023), which indicates that the standard Cobb-Douglas specification underestimates the mismatch index due to the absence of time-varying matching efficiency and elasticity.
-The steady increase in the mismatch index across job categories points to a growing dissonance between the skills offered by job seekers and the skills demanded by employers.
+The mismatch index across occupations increases to 0.6, which is significantly larger than Shibata (2020) and Higashi and Sasaki (2023), which indicates that the standard Cobb-Douglas specification underestimates the mismatch index due to the absence of time-varying matching efficiency and elasticity.
+The steady increase in the mismatch index across occupations points to a growing dissonance between the skills offered by job seekers and the skills demanded by employers.
 This trend may reflect evolving industry demands, technological advancements, and the lag in the workforce’s adaptation to new skills.
@@
 Finally, I extend the mismatch index proposed by Şahin et al.
-(2014) to the nonparametric version and find that the mismatch across job categories is more severe than across prefectures and the original Cobb-Douglas mismatch index is biased.
+(2014) to the nonparametric version and find that the mismatch across occupations is more severe than across prefectures and the original Cobb-Douglas mismatch index is biased.
 Acknowledgments

### v3

@@
 Blank cells mean no reported results.
-Second, this paper contributes to the literature on mismatch (Şahin et al.
-2014).
+Second, this paper contributes to the literature on labor market mismatch, building on the foundational work of Şahin et al.
+(2014) extending the earlier analysis by Jackman and Roper (1987).
+In the static matching model with multiple sectors, Jackman and Roper (1987) demonstrate that aggregate hires are maximized when unemployment is distributed across sectors in a way that equalizes sectoral labor-market tightness.
+Şahin et al.
+(2014) extend this framework by incorporating a dynamic, stochastic environment with multiple sources of heterogeneity, providing a methodology to construct counterfactual measures of unemployment in the absence of mismatch.
+Their focus is on mismatch unemployment, defined as unemployed workers searching in the “wrong” sector.
+[^2] Note that a distinct body of literature, such as Eeckhout and Kircher (2011), uses the term “mismatch” to describe employed individuals in suboptimal jobs, where the alignment between worker skills and firm capital is not ideal.
+Eeckhout (2018) provides an overview of this literature, identifying three primary causes of mismatch: search frictions (Eeckhout and Kircher 2011), stochastic sorting (Chade and Eeckhout 2016), and multidimensional types (Lindenlaub 2017).
+In addition, Guvenen et al.
+(2020) explores skill mismatch, which quantifies the gap between the skills required by an occupation and the abilities a worker possesses.
+Another approach, as in Shimer (2012), takes a flow perspective by decomposing unemployment volatility into transition rates between four labor force states: (i) permanent-contract employment, (ii) temporary-contract employment, (iii) unemployment, and (iv) non-participation, providing a foundation for calculating mismatch measures based on employment status.
 The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al.
@@
 Note that the formulation allows us to incorporate a rich set of additional equilibrium constraints and elements such as the government’s costs of moving unemployed workers from some markets to other markets for calibration studies, which is out of the scope of this paper.
-4 Monte Carlo simulation
-  • How large is the bias of matching efficiency from the Cobb Douglass specification?
-  • How large is the bias of mismatch from the Cobb Douglass specification?
-  • Mismatch literature:
-    • mismatch Shimer (2007)
-    • mismatch and large firms (Eeckhout 2018)
-    • Classical Jackman and Roper (1987)
 4 Data
@@
 (2014) to the nonparametric version and find that the mismatch across occupations is more severe than across prefectures and the original Cobb-Douglas mismatch index is biased.
-Acknowledgments
-I thank Fuhito Kojima, Kosuke Uetake, Akira Matsushita, Kazuhiro Teramoto, Ryo Kambayashi, Higashi Yudai for their valuable advice. This work was supported by JST ERATO Grant Number JPMJER2301, Japan.
 Appendix A Appendix

### v4

@@
 I find that the mismatch across occupations is more severe than across prefectures and the original Cobb-Douglas mismatch index is underestimated.
-Keywords: matching efficiency, mismatch
+Keywords: matching efficiency, matching elasticity, matching function, mismatch
 JEL code: E24, J61, J62, J64
@@
 [^1] Brancaccio et al.
-(2020a, 2023) apply the method to estimate matching function in exporter-ship transactions in the global bulk shipping market.
+(2020a, 2023) apply the method to estimate the matching function in exporter-ship transactions in the global bulk shipping market.
 Brancaccio et al.
@@
 Here, $H/U$ and $V/U$ are known as a job-finding rate and market tightness.
+It is known that increasing returns to scale in the matching process can create multiple equilibria and the subsequent literature has found scant evidence for increasing returns, as in (Petrongolo and Pissarides 2001).
 Then, by applying nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) proves that $G(H,U,V)$ identifies $F(A,U)$ and $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point denoted as $A_{0}$ of the support of $(A,U,V)$.
@@
 I construct a nonparametric version of a mismatch index which is a measure of the fraction of hires foregone due to mismatch.
-Suppose there are $G$ local markets in time $t$ with a given number of vacancies, unemployed workers, and aggregate matching efficiencies in the economy.
-The number of hires $H_{gt}$ in market $g$ in time $t$ is determined by a matching function $m(A_{gt}U_{gt},V_{gt})$.
+Suppose there are $L$ local markets in time $t$ with a given number of vacancies, unemployed workers, and aggregate matching efficiencies in the economy.
+The number of hires $H_{\ell t}$ in market $g$ in time $t$ is determined by a matching function $m(A_{\ell t}U_{\ell t},V_{\ell t})$.
 Note that, unlike Şahin et al.
-(2014), I allow any specification of $m$ satisfying CRS and independence of $A_{gt}$ and $U_{gt}$ conditional on $V_{gt}$.
-Given matching efficiency $A_{gt}$ and the number of vacancies $V_{gt}$, a social planner maximizes the total hires in the economy by distributing a given number of unemployed workers to each labor market.
+(2014), I allow any specification of $m$ satisfying CRS and independence of $A_{\ell t}$ and $U_{\ell t}$ conditional on $V_{\ell t}$.
+Given matching efficiency $A_{\ell t}$ and the number of vacancies $V_{\ell t}$, a social planner maximizes the total hires in the economy by distributing a given number of unemployed workers to each labor market.
 Following Şahin et al.
@@
 Under the assumptions of homogenous productivity and job separation rate across labor markets, the optimal allocation of unemployed workers satisfies the following equilibrium conditions:
-$\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)=\cdots=\frac{\partial m}{\partial U}\left(A_{Gt}U_{Gt}^{*},V_{Gt}\right)$ (1)
+$\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)=\cdots=\frac{\partial m}{\partial U}\left(A_{Lt}U_{Lt}^{*},V_{Lt}\right)$ (1)
 where * denotes the planner’s allocation.
-The equilibrium condition means that keeping total unemployed workers $U_{t}=\sum_{g=1}^{G}U_{gt}$, the planner allocates more unemployed workers to more effective labor markets, that is, with more vacancies and higher matching efficiency until their marginal contribution to the hires is equalized across markets.
+The equilibrium condition means that keeping total unemployed workers $U_{t}=\sum_{\ell=1}^{L}U_{\ell t}$, the planner allocates more unemployed workers to more effective labor markets, that is, with more vacancies and higher matching efficiency until their marginal contribution to the hires is equalized across markets.
 The main difference from Şahin et al.
@@
 In Section 3.2, I explain how to compute the optimal allocation of unemployed workers numerically.
-Using optimal allocation of unemployed workers $U_{gt}^{*}$, the aggregate actual and optimal number of new hires can be expressed as
-$H_{t}$ $=\sum_{g=1}^{G}m(A_{gt}U_{gt},V_{gt}),$
-$H_{t}^{*}$ $=\sum_{g=1}^{G}m(A_{gt}U_{gt}^{*},V_{gt}).$
+Using optimal allocation of unemployed workers $U_{\ell t}^{*}$, the aggregate actual and optimal number of new hires can be expressed as
+$H_{t}$ $=\sum_{\ell=1}^{L}m(A_{\ell t}U_{\ell t},V_{\ell t}),$
+$H_{t}^{*}$ $=\sum_{\ell=1}^{L}m(A_{\ell t}U_{\ell t}^{*},V_{\ell t}).$
 Using these expressions, I define mismatch index as
@@
 The constrained optimization problem in time $t$ is defined as follows.
-$\max_{U_{1t},\cdots,U_{Gt}}$ $\sum_{g=1}^{G}m_{gt}\left(A_{gt}U_{gt},V_{gt}\right),$ (3)
-subject to $\text{subject to }\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)$ $=\cdots=\frac{\partial m}{\partial U}\left(A_{Gt}U_{Gt}^{*},V_{Gt}\right),$
-$U_{t}$ $=\sum_{g=1}^{G}U_{gt}.$
+$\max_{U_{1t},\cdots,U_{Lt}}$ $\sum_{\ell=1}^{L}m\left(A_{\ell t}U_{\ell t},V_{\ell t}\right),$ (3)
+subject to $\text{subject to }\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)$ $=\cdots=\frac{\partial m}{\partial U}\left(A_{Lt}U_{Lt}^{*},V_{Lt}\right),$
+$U_{t}$ $=\sum_{\ell=1}^{L}U_{\ell t}.$
 Maximizing the objective function is equivalent to solving equilibrium conditions (1) so that the central planner’s problem represents a nonlinear system of equations with linear equilibrium constraints.
@@
 Note that the formulation allows us to incorporate a rich set of additional equilibrium constraints and elements such as the government’s costs of moving unemployed workers from some markets to other markets for calibration studies, which is out of the scope of this paper.
+4 Simulation
+  • [TBA]
 5 Data
@@
 Hello Work is a government-operated institution in Japan that provides job seekers with employment counseling, job placement services, and vocational training, playing a critical role in Japan’s labor market.
-The data is often used as in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018) estimating the traditional Cobb-Douglas matching function.
+The data is often used as in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), Kambayashi (2013), and Higashi (2018) estimating the traditional Cobb-Douglas matching function.
 Using up-to-date Hello Work data, Kawata and Sato (2021) construct a simple framework to quantitatively measure the impacts of an economic shock of COVID-19 on unemployed workers’ welfare in their companion project.
@@
 6.1 Aggregate trends in 1972-2023
-Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($M/U$ and $M/V$) in aggregate labor markets.
+Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($H/U$ and $H/V$) in aggregate labor markets.
 [^9] Worker finding rates are also known as vacancy yields.
@@
 In particular, the significant decline after 2015 is remarkable.
-This seems due to an increase in matching opportunities outside of the government-operated platform, which is discussed in Otani (2024) using proprietary on-the-job search platform data.
+This seems due to an increase in matching opportunities outside of the government-operated platform, which is discussed in Otani (2024) using proprietary on-the-job search platform data and Kanayama and Otani (2024) using spot gig work platform data.
 The implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (range: 0.5-0.7) and Japanese studies such as Higashi (2018) (0.38 for 2000-2014 monthly), Kawata (2019) (0.48 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.56 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.6 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.8 for 1996-2001 prefecture-month-level).
@@
 Appendix A Appendix
-A.1 Year-level trends in 1966-2023
+A.1 Simulation results
+  • How large is the bias of matching efficiency from the Cobb Douglass specification?
+  • How large is the bias of mismatch from the Cobb Douglass specification?
+  • Mismatch literature:
+    • mismatch Shimer (2007)
+    • mismatch and large firms (Eeckhout 2018)
+    • Classical Jackman and Roper (1987)
+  • We want to check robustness to specification, stationarity, and endogeneity.
+    • Restricting CRS class
+    • robustness to violation of independence of $A$ and $U$ conditional on $V$
+  • We want to confirm how large enough the sample size is.
+  • We want to quantify the bias of the standard Cobb-Douglas specification
+A.2 Year-level trends in 1966-2023
 Figure 12: Year-level results 1966-2024